Circular No. 7541

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 73P/SCHWASSMANN-WACHMANN 3 Z. Sekanina, Jet Propulsion Laboratory, writes: “Preliminary analysis

of the astrometric data reported for components C, B, and E indicates that E is not a recently separated fragment; it broke off from C in mid-Dec. 1995, some 85 ± 7 days after perihelion, with a relative deceleration of 7.6 ± 0.4 units (of 10^{−5} the solar attraction) and with a submeter-per-second relative velocity. Scenarios with E splitting off from B are unlikely. The failure to detect E before the comet’s 1996 conjunction with the sun was caused by its subarcsec separation from C until late Mar. 1996, when the comet already was < 20^{o} from the sun. However, unless too faint, E should have been detected in high-resolution images taken in Aug.–Sept. 1996, when it was 8′′–10′′ from C. Recent offsets of component B from C are consistent with those reported in 1995–1996, suggesting that B separated from C on 1995 Nov. 11 ± 4 with a relative deceleration of nearly 8 units and a relative velocity of ∼ 1.7 m/s. Predicted separation distances and position angles for B and E relative to C (0^{h} TT; p.a. are, within 1^{o}, the same for both companions): 2000 Dec. 12, 580′′, 1600′′, 295^{o}; Dec. 22, 624′′, 1730′′, 292^{o}; 2001 Jan. 1, 656′′, 1830′′, 288^{o}; Apr. 21, 271′′, 760′′, 245^{o}; May 1, 240′′, 670′′, 244^{o}; May 11, 213′′, 590′′, 243^{o}; May 21, 191′′, 530′′, 242^{o}; May 31, 173′′, 470′′, 241^{o}.”

KS 1947+300 I. Negueruela, Observatoire de Strasbourg; A. Marco, University of

Alicante; R. Speziali and G. L. Israel, Osservatorio di Roma, write: “On Dec. 1, we obtained a low-resolution spectrum of the only blue star in the error circle of KS 1947+300 (IAUC 7523; Goranskij et al. 1991, Sov. Astron. Lett. 17, 938) with the 1.82-m telescope (+ AFOSC grism no. 4) at the Asiago Observatory. The object displays relatively strong Hα emission (EW = 1.5 nm) and weak emission in Hβ. The spectrum is typical of a moderately reddened OBe star in a Be/x-ray binary system, as expected from the identification with GRO J1948+32 (IAUC 7531). The spectral type is approximately B0. Infrared photometry was carried out on Dec. 6 with the AZT-24 1.1-m telescope (+ SWIRCAM) at Campo Imperatore, yielding the following magnitudes: J = 12.0 ± 0.2, H = 11.5 ± 0.1, K = 11.0 ± 0.2 (suggesting the presence of circumstellar emission). This is most likely the optical counterpart to KS 1947+300.”

2000 December 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green